# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v6

@@ V-C Plan Selection
V-D Periodic State Update
@@
V-E Multi-agent Reinforcement Learning
Our approach provides a generalized model for multi-agent reinforcement learning by adopting the framework of centralized training and decentralized execution, as shown in Fig.4. During execution, each agent obtains a local aggregated observation comprising both private and public information. Apart from their own private data such as locations, battery levels, and selected plans, this observation contains aggregated plans shared via the tree communication model (shown in Fig. 2). The training part is built based on an actor-critic model. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The critic, however, evaluates the actions of drones and guide them toward collectively optimal decisions. In addition, DO-RL employs Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process [37].
@@ V-E Multi-agent Reinforcement Learning
where $\gamma$ is a discount factor; $\hat{A}^{u}_{k}$ is an advantage function. Then, the critic network provides $\hat{A}^{u}_{k}$ to the actor network to increase the probability of actions that have a positive impact and decrease the ones that have negative impact. The actions are taken by drones as $a^{u}_{t}=\pi(\textbf{o}^{u}_{t})$. The parameter $\theta^{\pi}$ of the actor network is updated by maximizing the clip objective $L_{CLIP}(\theta^{\pi})$, as follows [37]:
@@ VI-A Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The number of vehicles serves as the required sensing value in the model. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^2] https://www.eclipse.org/sumo/ to accurately generate realistic flows of vehicles. Fig. 5 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 6 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations. Noted that the simulated traffic flow can be displayed in a big screen in a testbed experimentation using hardware to emulate the complex outdoor sensing environments, which has been proved by our previous work in [15, 34].
@@
Drones. We use DJI Phantom 4 Pro model [^3] https://www.dji.com/uk/phantom-4-pro/info. Each drone has a body weight of $1.07$ kg and four propellers, each with a diameter of $0.35$ m. The ground speed of the drones is set at $6.94$ m/s, and the drag force is determined to be $4.1134$[36]. As a consequence, the power consumption remains consistent across all drones. Additionally, drones are equipped with a $6000$ mAh LiPo 2S battery ($0.31$ kg). The maximum flying time for drones is approximately $30$ min, which is set as one flying period. A flying period has $S=30$ timeslots, each of one minute. Moreover, to ensure that the camera covers the entire area of a cell (see Fig. 5), the minimum hovering height of drones is determined at which the field of view of the camera and the cell overlap. The hovering height $h$ is calculated based on the distance between any two cells (approximately $200$ meters), the camera resolution, focal length derived from camera calibration, and ground sampling distance [38]. Therefore, each drone equipped with a 4K camera senses from a minimum height of $82.4$ meters.
@@
Planning and distributed optimization. We leverage the decentralized multi-agent collective learning method of EPOS [^4] EPOS is open-source and available at: https://github.com/epournaras/EPOS., the Economic Planning and Optimized Selections [7]. We generate $L=64$ plans for each agent. All generated plans are made openly available to encourage further research on coordinated spatio-temporal sensing of drones [15, 24, 12]. Each agent in EPOS is mapped to a drone. During the coordinated plan selection via EPOS, agents self-organize into a balanced binary tree as a way of structuring efficient learning interactions [39]. The shared goal of the agents is to minimize the RMSE between the total sensing values collected and the target via Eq.(7). In the one execution of EPOS, the agents perform $40$ bottom-up and top-down learning iterations during which RMSE converges to a minimum optimized value.
@@
Complexity. Given the number of nodes per layer $W$, number of episodes $\mathcal{E}$, size of state space $\mathcal{X}$, and number of actions $A$, the computational complexity of DNN is approximately $O(C_{dnn}):=O((\mathcal{X}+A)W+3W^{2})$ [25]. The comparison of both computational and communication cost is shown in Table II, where $I$ denotes the number of iterations in EPOS. The results illustrate that DO-RL significantly lowers the computation cost by reducing the number of actions required during training on a temporal scale, compared to the state-of-the-art MAPPO. DO-RL also has low communication cost due to the tree communication structure.
TABLE II: Comparison of computational and communication costs.
@@
EPOS [7]
MAPPO [16]
@@ VI-B Baselines and Metrics
Greedy. It aims to assign each drone to a single cell, enabling it to accomplish the necessary sensing tasks within a time period [20]. This algorithm minimizes the number of visited cells and travel distance, resulting in reduced energy consumption for drones. To promote decentralization, we implement the algorithm with a local view for each drone, i.e., they have no coordination or awareness of cell occupancy by other drones. By adopting this approach, we ensure an energy-efficient and independent operation of each drone.
EPOS. It is designed to select the optimal plan for drones based on the generated plans [7]. It incorporates the plan generation and the periodic state update, but does not support any long-term strategic navigation. Agents randomly choose any charging station to land on at every period.
MAPPO. It is one state-of-the-art DRL algorithm using PPO [16, 37], but does not include distributed optimization compared to DO-RL. Moreover, agents in MAPPO learn the flying directions timeslot by timeslot, rather than period by period as in DO-RL. At each timeslot, a drone takes actions to horizontally move to an adjacent cell in eight directions or hover (similar to the actions in DO-RL), and returns to the nearest charging station at the end of every $S$ timeslots (one period). For fair comparison, this method employs the same reward function and structured tree communication model to share the aggregated observation.
@@ VI-C Results and Analysis
Number of drones/cells with fixed drones density. If we fix the density of drones but increase the number of drones, the number of cells increases proportionally. For example, we use the density of drones in the basic scenario of $0.25$, and increase the number of drones from $9$ to $49$ and the number of cells from $36$ to $196$. As shown in Fig. 10, when the number of drones/cells increases, the sensing accuracy of all methods increases. However, the mission efficiency of both EPOS and MAPPO decreases (by $35.93\%$ and $26.42\%$ respectively). This is due to EPOS facing challenges in efficiently scheduling drones for a large number of cells, while MAPPO experiences increased computational complexity. In contrast, DO-RL overcomes these challenges and demonstrates a $4.62\%$ increase in mission efficiency. When increasing to $49$ drones and $196$ cells, however, the sensing accuracy of DO-RL degrades, decreased by $15.96\%$, due to the high complexity of scheduling and computation. In overall, DO-RL has statistically higher overall performance than other methods, with an average improvement of $46.98\%$ (maximum p-value less than $0.001$).
@@
VII Conclusion and Future Work
This paper studies a large-scale spatio-temporal task assignment problem by a swarm of autonomous drones. We propose a new synergetic optimization approach named DO-RL, by integrating both evolutionary distributed optimization and DRL, to solve the problem by optimizing the flight paths, data collection and recharging assignment of drone swarms. In each time period, drones determine their overall flying directions to destinations with charging stations via DRL, and autonomously evolve their navigation and sensing plans using collective learning. Extensive experimentation using realistic urban mobility reveals the effectiveness of DO-RL and provides valuable insights for the broader community: The combination of short-term and long-term strategies proves highly effective in overcoming their standalone limitations in drone-based traffic monitoring. Short-term methods compensate the challenges of training complexity, energy consumption and recharging, while long-term approaches account for maximizing accumulated rewards, enhancing the overall sensing performance. This collaborative approach allows the system to dynamically adapt to changing environments, efficiently managing varying drone resources and enabling extended operations.
@@ Appendix A Mathematical Notations
Table III illustrates the list of mathematical notations used in this paper.
TABLE III: Notations.
@@ 8                   Pick a plan $P^{u}(a^{u},t)$ and share it with other drones via tree communication;
9                   Observe the aggregated plan $P^{-u}(a^{u},t)$ and pick an optimized plan through collective learning [7];
@@ Appendix C Effect of different parameters
Agents’ behavior. Fig. 13 illustrates the effect of agents’ behavior by varying the parameter of $\beta$ [7]. As $\beta$ increases from 00 to $1$, agents reduce the energy cost of their selected plans, while decreasing both efficiency and accuracy. This is because, according to Eq. 7, drones with higher $\beta$ choose a plan with lower energy cost, which degrades the matching between the total sensing and the target. Since the overall performance for $\beta\in[0.1,0.8]$ is statistically similar, using the Mann-Whitney U test with average p-value of $0.0003$, drones are allowed to choose their behaviors within this range.
@@ Appendix D Power consumption model
Based on the model in [36], the power consumption with forward velocity and forward pitch is given by:

### v7

@@ 5.3 Plan Selection
5.4 Periodic State Update
@@
5.5 Multi-agent Reinforcement Learning
Our approach provides a generalized model for multi-agent reinforcement learning by adopting the framework of centralized training and decentralized execution, as shown in Fig.4. During execution, each agent obtains a local aggregated observation comprising both private and public information. Apart from their own private data such as locations, battery levels, and selected plans, this observation contains aggregated plans shared via the tree communication model (shown in Fig. 2). The training part is built based on an actor-critic model. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The critic, however, evaluates the actions of drones and guide them toward collectively optimal decisions. In addition, DO-RL employs Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process Yi et al. (2022).
@@ 5.5 Multi-agent Reinforcement Learning
where $\gamma$ is a discount factor; $\hat{A}^{u}_{k}$ is an advantage function. Then, the critic network provides $\hat{A}^{u}_{k}$ to the actor network to increase the probability of actions that have a positive impact and decrease the ones that have negative impact. The actions are taken by drones as $a^{u}_{t}=\pi(\textbf{o}^{u}_{t})$. The parameter $\theta^{\pi}$ of the actor network is updated by maximizing the clip objective $L_{CLIP}(\theta^{\pi})$, as follows Yi et al. (2022):
@@ 6.1 Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The number of vehicles serves as the required sensing value in the model. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^3] https://www.eclipse.org/sumo/ to accurately generate realistic flows of vehicles. Fig. 5 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 6 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations.
@@
Drones. We use DJI Phantom 4 Pro model [^4] https://www.dji.com/uk/phantom-4-pro/info. Each drone has a body weight of $1.07$ kg and four propellers, each with a diameter of $0.35$ m. The ground speed of the drones is set at $6.94$ m/s, and the drag force is determined to be $4.1134$Monwar et al. (2018). As a consequence, the power consumption remains consistent across all drones. Additionally, drones are equipped with a $6000$ mAh LiPo 2S battery ($0.31$ kg). The maximum flying time for drones is approximately $30$ min, which is set as one flying period. A flying period has $S=30$ timeslots, each of one minute. Moreover, to ensure that the camera covers the entire area of a cell (see Fig. 5), the minimum hovering height of drones is determined at which the field of view of the camera and the cell overlap. The hovering height $h$ is calculated based on the distance between any two cells (approximately $200$ meters), the camera resolution, focal length derived from camera calibration, and ground sampling distance Wierzbicki (2018). Therefore, each drone equipped with a 4K camera senses from a minimum height of $82.4$ meters.
@@
Planning and distributed optimization. We leverage the decentralized multi-agent collective learning method of EPOS [^5] EPOS is open-source and available at: https://github.com/epournaras/EPOS., the Economic Planning and Optimized Selections Pournaras et al. (2018). We generate $L=64$ plans for each agent. All generated plans are made openly available to encourage further research on coordinated spatio-temporal sensing of drones Qin and Pournaras (2023). Each agent in EPOS is mapped to a drone. During the coordinated plan selection via EPOS, agents self-organize into a balanced binary tree as a way of structuring efficient learning interactions Pournaras (2013). The shared goal of the agents is to minimize the RMSE between the total sensing values collected and the target via Eq.(7). In the one execution of EPOS, the agents perform $40$ bottom-up and top-down learning iterations during which RMSE converges to a minimum optimized value.
@@
Complexity. Given the number of nodes per layer $W$, number of episodes $\mathcal{E}$, size of state space $\mathcal{X}$, and number of actions $A$, the computational complexity of DNN is approximately $O(C_{dnn}):=O((\mathcal{X}+A)W+3W^{2})$ Omoniwa et al. (2023). The comparison of both computational and communication cost is shown in Table 2, where $I$ denotes the number of iterations in EPOS. The results illustrate that DO-RL significantly lowers the computation cost by reducing the number of actions required during training on a temporal scale, compared to the state-of-the-art MAPPO. DO-RL also has low communication cost due to the tree communication structure.
Table 2: Comparison of computational and communication costs.
@@
EPOS Pournaras et al. (2018)
MAPPO Ding et al. (2021)
@@ 6.2 Baselines and Metrics
Greedy. It aims to assign each drone to a single cell, enabling it to accomplish the necessary sensing tasks within a time period Bartolini et al. (2019). This algorithm minimizes the number of visited cells and travel distance, resulting in reduced energy consumption for drones. To promote decentralization, we implement the algorithm with a local view for each drone, i.e., they have no coordination or awareness of cell occupancy by other drones. By adopting this approach, we ensure an energy-efficient and independent operation of each drone.
EPOS. It is designed to select the optimal plan for drones based on the generated plans Pournaras et al. (2018). It incorporates the plan generation and the periodic state update, but does not support any long-term strategic navigation. Agents randomly choose any charging station to land on at every period.
MAPPO. It is one state-of-the-art DRL algorithm using PPO Ding et al. (2021); Yi et al. (2022), but does not include distributed optimization compared to DO-RL. Moreover, agents in MAPPO learn the flying directions timeslot by timeslot, rather than period by period as in DO-RL. At each timeslot, a drone takes actions to horizontally move to an adjacent cell in eight directions or hover (similar to the actions in DO-RL), and returns to the nearest charging station at the end of every $S$ timeslots (one period). For fair comparison, this method employs the same reward function and structured tree communication model to share the aggregated observation.
@@ 6.3 Results and Analysis
Number of drones/cells with fixed drones density. If we fix the density of drones but increase the number of drones, the number of cells increases proportionally. For example, we use the density of drones in the basic scenario of $0.25$, and increase the number of drones from $9$ to $49$ and the number of cells from $36$ to $196$. As shown in Fig. 10, when the number of drones/cells increases, the sensing accuracy of all methods increases. However, the mission efficiency of both EPOS and MAPPO decreases (by $35.93\%$ and $26.42\%$ respectively). This is due to EPOS facing challenges in efficiently scheduling drones for a large number of cells, while MAPPO experiences increased computational complexity. In contrast, DO-RL overcomes these challenges and demonstrates a $4.62\%$ increase in mission efficiency. When increasing to $49$ drones and $196$ cells, however, the sensing accuracy of DO-RL degrades, decreased by $15.96\%$, due to the high complexity of scheduling and computation. In overall, DO-RL has statistically higher overall performance than other methods, with an average improvement of $46.98\%$ (maximum p-value less than $0.001$), which proves the efficiency of the proposed approach.
@@
Apart from sensing performance, we study battery consumption and charging of DO-RL in complex scenario.
@@
7 Conclusion and Future Work
This paper studies a large-scale spatio-temporal task assignment problem by a swarm of autonomous drones. We propose a new synergetic optimization approach named DO-RL, by integrating both distributed evolutionary optimization and DRL, to solve the problem by optimizing the flight paths, data collection and recharging assignment of drone swarms. In each time period, drones determine their overall flying directions to destinations with charging stations via DRL, and autonomously evolve their navigation and sensing plans using collective learning. Extensive experimentation using realistic urban mobility reveals the effectiveness of DO-RL and provides valuable insights for the broader community: The combination of short-term and long-term strategies proves highly effective in overcoming their standalone limitations in drone-based traffic monitoring. Short-term methods compensate the challenges of training complexity, energy consumption and recharging, while long-term approaches account for maximizing accumulated rewards, enhancing the overall sensing performance. This collaborative approach allows the system to dynamically adapt to changing environments, efficiently managing varying drone resources and enabling extended operations.
@@ Appendix A Mathematical Notations
Table 3 illustrates the list of mathematical notations used in this paper.
Table 3: Notations.
@@ 8                   Pick a plan $P^{u}(a^{u},t)$ and share it with other drones via tree communication;
9                   Observe the aggregated plan $P^{-u}(a^{u},t)$ and pick an optimized plan through collective learning Pournaras et al. (2018);
@@ Appendix C Effect of different parameters
Agents’ behavior. Fig. 13 illustrates the effect of agents’ behavior by varying the parameter of $\beta$ Pournaras et al. (2018). As $\beta$ increases from 00 to $1$, agents reduce the energy cost of their selected plans, while decreasing both efficiency and accuracy. This is because, according to Eq. 7, drones with higher $\beta$ choose a plan with lower energy cost, which degrades the matching between the total sensing and the target. Since the overall performance for $\beta\in[0.1,0.8]$ is statistically similar, using the Mann-Whitney U test with average p-value of $0.0003$, drones are allowed to choose their behaviors within this range.
@@ Appendix D Power consumption model
Based on the model in Monwar et al. (2018), the power consumption with forward velocity and forward pitch is given by: